OpenStudy (anonymous):

Jake and Amanda drove for 9 hours against a wind speed of 8 mph. Traveling with the same wind speed, it took them 7 hours to make the return trip. How fast were they traveling in still air?

OpenStudy (anonymous):

This can be answered with a proportion. \[\frac{ 8 }{ 9 } = \frac{x}{7}\] Then cross multiply. 9x=56 \[\frac{56}{9} = 6\frac{2}{9}\]